Assume that gear a has 56 teeth and gear b has 14 teeth. Suppose that the number of counterclockwise rotations is designated by a positive number and the number of clockwise rotations by a negative number. If gear A rotates 7 times per minute, how many times per minute does gear b rotate? Explain
Answer by dfrazzetto(283)

Ratio of gears:
A/B = 56/14 = 4, so for every 1 revolution gear A makes, gear b will make 4.
Gear A's rotational speed is given as positive, so we can assume it is counterclockwise and gear B is clockwise then:
RPM of B = -4(RPM of A) = -4(7) = -28 rotations per minute